# Advancing large-scale thin-film PPLN nonlinear photonics with segmented tunable micro-heaters

### 15 Dec 2023 v1

@@ IV CONCLUSIONS
In conclusion, we have demonstrated the wafer-scale production of TF-PPLN optical waveguides leveraging UV stepper lithography and an automated poling probe station. We address the degradation of conversion efficiency due to inhomogeneous film thickness by employing a segmented thermal tuning scheme. We show the successful recovery of single-peak QPM spectral profiles with up to 108 enhancement of the peak conversion efficiency without the need of pre-fabrication thickness mapping or design compensation, which is highly appealing for high-volume and low-cost wafer-scale production. The thermal tuning efficiency can be further enhanced by incorporating local air trenches to minimize heat leakage [26]. Even higher peak conversion efficiencies and better QPM spectral shapes could be achieved by implementing more thermal tuning modules and an automated control algorithm for optimizing the tuning parameters. This will enable faster searching for optimal working points, simultaneous control over multiple TF-PPLN devices, and real-time adaptation to environmental drifts. The scalable fabrication and tuning methodologies presented in this work mark an important step towards future large-scale nonlinear photonic integrated circuits with high efficiencies, versatile functionalities, and excellent reconfigurability, unlocking new opportunities for future quantum and classical photonic applications.
Acknowledgements.
The author would like to thank Dr. Wing-Han Wong, Mr. Hanke FENG, and Dr. Baojie CHEN for his kind help and guidance in the fabrication process. This work is supported by Research Grants Council, University Grants Committee (CityU 11204820, NCityU113/20); Croucher Foundation (9509005). Xiaoting Li acknowledge support form Hong Kong PhD Fellowship Scheme (PF18-17958). The authors declare no conflicts of interest.References
- Boes et al. [2023] A. Boes, L. Chang, C. Langrock, M. Yu, M. Zhang, Q. Lin, M. Lončar, M. Fejer, J. Bowers, and A. Mitchell, Lithium niobate photonics: Unlocking the electromagnetic spectrum, Science 379, eabj4396 (2023).
- Wang et al. [2018a] C. Wang, C. Langrock, A. Marandi, M. Jankowski, M. Zhang, B. Desiatov, M. M. Fejer, and M. Lončar, Ultrahigh-efficiency wavelength conversion in nanophotonic periodically poled lithium niobate waveguides, Optica 5, 1438 (2018a).
- Zhao et al. [2020a] J. Zhao, M. Rüsing, U. A. Javid, J. Ling, M. Li, Q. Lin, and S. Mookherjea, Shallow-etched thin-film lithium niobate waveguides for highly-efficient second-harmonic generation, Optics Express 28, 19669 (2020a).
- Rao et al. [2019] A. Rao, K. Abdelsalam, T. Sjaardema, A. Honardoost, G. F. Camacho-Gonzalez, and S. Fathpour, Actively-monitored periodic-poling in thin-film lithium niobate photonic waveguides with ultrahigh nonlinear conversion efficiency of 4600 %W -1 cm -2, Optics Express 27, 25920 (2019).
- Lu et al. [2019] J. Lu, J. B. Surya, X. Liu, A. W. Bruch, Z. Gong, Y. Xu, and H. X. Tang, Periodically poled thin-film lithium niobate microring resonators with a second-harmonic generation efficiency of 250,000%/W, Optica 6, 1455 (2019).
- Chen et al. [2019] J.-Y. Chen, Z.-H. Ma, Y. M. Sua, Z. Li, C. Tang, and Y.-P. Huang, Ultra-efficient frequency conversion in quasi-phase-matched lithium niobate microrings, Optica 6, 1244 (2019).
- Ledezma et al. [2022] L. Ledezma, R. Sekine, Q. Guo, R. Nehra, S. Jahani, and A. Marandi, Intense optical parametric amplification in dispersion-engineered nanophotonic lithium niobate waveguides, Optica 9, 303 (2022).
- Jankowski et al. [2022] M. Jankowski, N. Jornod, C. Langrock, B. Desiatov, A. Marandi, M. Lončar, and M. M. Fejer, Quasi-static optical parametric amplification, Optica 9, 273 (2022).
- Zhao et al. [2020b] J. Zhao, C. Ma, M. Rüsing, and S. Mookherjea, High Quality Entangled Photon Pair Generation in Periodically Poled Thin-Film Lithium Niobate Waveguides, Physical Review Letters 124, 163603 (2020b).
- Xue et al. [2021a] G.-T. Xue, X.-H. Tian, C. Zhang, Z. Xie, P. Xu, Y.-X. Gong, and S.-N. Zhu, Effect of thickness variations of lithium niobate on insulator waveguide on the frequency spectrum of spontaneous parametric down-conversion*, Chinese Physics B 30, 110313 (2021a).
- Wang et al. [2018b] C. Wang, M. Zhang, X. Chen, M. Bertrand, A. Shams-Ansari, S. Chandrasekhar, P. Winzer, and M. Lončar, Integrated lithium niobate electro-optic modulators operating at CMOS-compatible voltages, Nature 562, 101 (2018b).
- He et al. [2019a] M. He, M. Xu, Y. Ren, J. Jian, Z. Ruan, Y. Xu, S. Gao, S. Sun, X. Wen, L. Zhou, L. Liu, C. Guo, H. Chen, S. Yu, L. Liu, and X. Cai, High-performance hybrid silicon and lithium niobate Mach–Zehnder modulators for 100 Gbit s-1 and beyond, Nature Photonics 13, 359 (2019a).
- Wan et al. [2022] L. Wan, Z. Yang, W. Zhou, M. Wen, T. Feng, S. Zeng, D. Liu, H. Li, J. Pan, N. Zhu, W. Liu, and Z. Li, Highly efficient acousto-optic modulation using nonsuspended thin-film lithium niobate-chalcogenide hybrid waveguides, Light: Science & Applications 11, 145 (2022).
- Wang et al. [2019] C. Wang, M. Zhang, M. Yu, R. Zhu, H. Hu, and M. Loncar, Monolithic lithium niobate photonic circuits for Kerr frequency comb generation and modulation, Nature Communications 10, 978 (2019).
- He et al. [2019b] Y. He, Q.-F. Yang, J. Ling, R. Luo, H. Liang, M. Li, B. Shen, H. Wang, K. Vahala, and Q. Lin, Self-starting bi-chromatic LiNbO 3 soliton microcomb, Optica 6, 1138 (2019b).
- Bruch et al. [2021] A. W. Bruch, X. Liu, Z. Gong, J. B. Surya, M. Li, C.-L. Zou, and H. X. Tang, Pockels soliton microcomb, Nature Photonics 15, 21 (2021).
- Op De Beeck et al. [2021] C. Op De Beeck, F. M. Mayor, S. Cuyvers, S. Poelman, J. F. Herrmann, O. Atalar, T. P. McKenna, B. Haq, W. Jiang, J. D. Witmer, G. Roelkens, A. H. Safavi-Naeini, R. Van Laer, and B. Kuyken, III/V-on-lithium niobate amplifiers and lasers, Optica 8, 1288 (2021).
- Desiatov and Lončar [2019] B. Desiatov and M. Lončar, Silicon photodetector for integrated lithium niobate photonics, Applied Physics Letters 115, 121108 (2019).
- Guo et al. [2022a] X. Guo, L. Shao, L. He, K. Luke, J. Morgan, K. Sun, J. Gao, T.-C. Tzu, Y. Shen, D. Chen, B. Guo, F. Yu, Q. Yu, M. Jafari, M. Lončar, M. Zhang, and A. Beling, High-performance modified uni-traveling carrier photodiode integrated on a thin-film lithium niobate platform, Photonics Research 10, 1338 (2022a).
- Zhu et al. [2023] S. Zhu, Y. Zhang, Y. Ren, Y. Wang, K. Zhai, H. Feng, Y. Jin, Z. Lin, J. Feng, S. Li, Q. Yang, N. H. Zhu, E. Y.-B. Pun, and C. Wang, Waveguide-Integrated Two-Dimensional Material Photodetectors in Thin-Film Lithium Niobate, Advanced Photonics Research 4, 2300045 (2023).
- Nehra et al. [2022] R. Nehra, R. Sekine, L. Ledezma, Q. Guo, R. M. Gray, A. Roy, and A. Marandi, Few-cycle vacuum squeezing in nanophotonics, Science 377, 1333 (2022).
- Stokowski et al. [2023] H. S. Stokowski, T. P. McKenna, T. Park, A. Y. Hwang, D. J. Dean, O. T. Celik, V. Ansari, M. M. Fejer, and A. H. Safavi-Naeini, Integrated quantum optical phase sensor in thin film lithium niobate, Nature Communications 14, 3355 (2023).
- Guo et al. [2022b] Q. Guo, R. Sekine, L. Ledezma, R. Nehra, D. J. Dean, A. Roy, R. M. Gray, S. Jahani, and A. Marandi, Femtojoule femtosecond all-optical switching in lithium niobate nanophotonics, Nature Photonics 16, 625 (2022b).
- Ledezma et al. [2023] L. Ledezma, A. Roy, L. Costa, R. Sekine, R. Gray, Q. Guo, R. Nehra, R. M. Briggs, and A. Marandi, Octave-spanning tunable infrared parametric oscillators in nanophotonics, Science Advances 9, eadf9711 (2023).
- Li et al. [2022] M. Li, L. Chang, L. Wu, J. Staffa, J. Ling, U. A. Javid, S. Xue, Y. He, R. Lopez-rios, T. J. Morin, H. Wang, B. Shen, S. Zeng, L. Zhu, K. J. Vahala, J. E. Bowers, and Q. Lin, Integrated Pockels laser, Nature Communications 13, 5344 (2022).
- Liu et al. [2022] X. Liu, C. Zhang, Y. Pan, R. Ma, X. Zhang, M. Chen, L. Liu, Z. Xie, S. Zhu, S. Yu, and X. Cai, Thermally tunable and efficient second-harmonic generation on thin-film lithium niobate with integrated micro-heater, Optics Letters 47, 4921 (2022).
- Luke et al. [2020] K. Luke, P. Kharel, C. Reimer, L. He, M. Loncar, and M. Zhang, Wafer-scale low-loss lithium niobate photonic integrated circuits, Optics Express 28, 24452 (2020).
- Fejer et al. [1992] M. Fejer, G. Magel, D. Jundt, and R. Byer, Quasi-phase-matched second harmonic generation: Tuning and tolerances, IEEE Journal of Quantum Electronics 28, 2631 (Nov./1992).
- Tian et al. [2021] X.-H. Tian, W. Zhou, K.-Q. Ren, C. Zhang, X. Liu, G.-T. Xue, J.-C. Duan, X. Cai, X. Hu, Y.-X. Gong, Z. Xie, and S.-N. Zhu, Effect of dimension variation for second-harmonic generation in lithium niobate on insulator waveguide [Invited], Chinese Optics Letters 19, 060015 (2021).
- Xue et al. [2021b] G.-T. Xue, Y.-F. Niu, X. Liu, J.-C. Duan, W. Chen, Y. Pan, K. Jia, X. Wang, H.-Y. Liu, Y. Zhang, P. Xu, G. Zhao, X. Cai, Y.-X. Gong, X. Hu, Z. Xie, and S. Zhu, Ultrabright Multiplexed Energy-Time-Entangled Photon Generation from Lithium Niobate on Insulator Chip, Physical Review Applied 15, 064059 (2021b).
- Santandrea et al. [2019] M. Santandrea, M. Stefszky, G. Roeland, and C. Silberhorn, Characterisation of fabrication inhomogeneities in Ti:LiNbO 3 waveguides, New Journal of Physics 21, 123005 (2019).
- Zhao et al. [2023] J. Zhao, X. Li, T.-C. Hu, A. A. Sayem, H. Li, A. Tate, K. Kim, R. Kopf, P. Sanjari, M. Earnshaw, N. K. Fontaine, C. Wang, and A. Blanco-Redondo, Unveiling the origins of quasi-phase matching spectral imperfections in thin-film lithium niobate frequency doublers, APL Photonics 8, 126106 (2023).
- Kuo [2022] P. S. Kuo, Noncritical phasematching behavior in thin-film lithium niobate frequency converters, Optics Letters 47, 54 (2022).
- Chen et al. [2023] P.-K. Chen, I. Briggs, C. Cui, L. Zhang, M. Shah, and L. Fan, Adapted poling to break the nonlinear efficiency limit in nanophotonic lithium niobate waveguides, Nature Nanotechnology 10.1038/s41565-023-01525-w (2023).
- Zhao et al. [2019] J. Zhao, M. Rüsing, and S. Mookherjea, Optical diagnostic methods for monitoring the poling of thin-film lithium niobate waveguides, Optics Express 27, 12025 (2019).
- Nagy and Reano [2020] J. T. Nagy and R. M. Reano, Submicrometer periodic poling of lithium niobate thin films with bipolar preconditioning pulses, Optical Materials Express 10, 1911 (2020).
- Niu et al. [2020] Y. Niu, C. Lin, X. Liu, Y. Chen, X. Hu, Y. Zhang, X. Cai, Y.-X. Gong, Z. Xie, and S. Zhu, Optimizing the efficiency of a periodically poled LNOI waveguide using in situ monitoring of the ferroelectric domains, Applied Physics Letters 116, 101104 (2020).

### 15 Mar 2024 v2

@@ 4 CONCLUSIONS AND DISCUSSIONS
In conclusion, we have demonstrated the wafer-scale production of TF-PPLN optical waveguides leveraging UV stepper lithography and an automated poling probe station. We address the degradation of conversion efficiency due to inhomogeneous film thickness by employing a segmented thermal tuning scheme. We show the successful recovery of single-peak QPM spectral profiles with up to 57 enhancement of the peak conversion efficiency, and achieve a highest normalized conversion efficiency of 3802Wcm in a 6 mm long device. Importantly, this is achieved without the need of pre-fabrication thickness mapping or design compensation, which is highly appealing for high-volume and low-cost wafer-scale production. The thermal tuning efficiency can be further enhanced by incorporating local air trenches to minimize heat leakage [26]. LNOI wafers with better initial thickness variations are also expected to reduce the required heating powers in our devices. Even higher peak conversion efficiencies and better QPM spectral shapes could be achieved by implementing more thermal tuning modules and an automated control algorithm for optimizing the tuning parameters. This will enable faster searching for optimal working points, simultaneous control over multiple TF-PPLN devices, and real-time adaptation to environmental drifts. Our segmented heater design could also be combined with the adaptive poling method to compensate for the remaining inhomogeneous broadening effects and facilitate active tuning of QPM wavelengths. The scalable fabrication and tuning methodologies presented in this work mark an important step towards future large-scale nonlinear photonic integrated circuits with high efficiencies, versatile functionalities, and excellent reconfigurability, unlocking new opportunities for future quantum and classical photonic applications.
Funding Research Grants Council, University Grants Committee (CityU 11204820, NCityU113/20); Croucher Foundation (9509005); Hong Kong PhD Fellowship Scheme (PF18-17958); National Natural Science Foundation of China (62105374).
Acknowledgments The author would like to thank Dr. Wing-Han Wong and Mr. Hanke FENG for their kind help and guidance in the fabrication process.
Disclosures The authors declare no conflicts of interest.
Data availability Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.